Determine the symmetry operation product for this equation.
C2 x C3^2 x S4 x C4 x i x σh x C3 x σ^4 x E x S2 x C2 =
Answer #1

Let us do it part by part.

C​​​​​​3​​​​​​​​​​2 × C​3 = E

E × X = X ( X is any arbitrary operation )

\sigma 4 = E ( even number of sigma planes leaves the coordinates unchanged )

C​​​​​​2 × C​​​​​​2 = E

S​​​​​​2 × i = E

C​​​​​​4 × \sigma h = S​​​​​​4

S​​​​​​4 ​​​​​× S 4 = E

So from this point of view we see that the operations gets reduced to the identity operation E.